# Chiral Floquet Engineering on Topological Fermions in Chiral Crystals

### v1

@@ Abstract
Fig. 1 $\bm{|}$: Multiple types of topological fermions and chiral Floquet engineering in the CoSi compound. a A schema of topological fermions at $\Gamma$ and R points which exhibit spin-1 (spin-3/2) and double Weyl (double spin-1) excitations without (with) SOC around the Fermi level in CoSi. b Schema of the left-circularly polarized light (LCPL)- and right-CPL (RCPL)- induced momentum shifts for topological fermions, marked as blue and red arrows, in the LH and RH CoSi crystal via Floquet engineering. The black cube represents the Brillouin zone (BZ) of the CoSi family and the red negative and blue positive signs denote the sign of topological charges.
@@
In this Letter, we report the chiral Floquet engineering on topological fermions in the CoSi family using the Floquet effective $\mathbf{k\cdot p}$ model from the perturbation theory, complemented by the Floquet tight-binding Hamiltonian based on the ab initio calculations as a benchmark. We show that under CPL pumping, gapless topological fermions with different topological charges undergo momentum shifts along opposite directions (see Fig. 1b) while preserving their topological properties. This phenomenon is distinct from the momentum shifts of Weyl points driven by the light-induced shear phonon mode in WTe2[55] and shifts of topological fermions under the linearly polarized light (LPL)[56]. The sign and magnitude for light-induced momentum shifts at $\Gamma$ and R points are determined by the Floquet chirality index $\Xi_{k=\Gamma,\rm{R}}$, which depends on the light-matter coupling strength and the interplay among the three distinct chiralities, the crystal handedness $\eta$, the chirality $\chi_{k=\Gamma,\rm{R}}(\eta)$ of topological fermions, and the chirality $\gamma$ of CPL. Via analyzing the Lie algebra representations of $\mathfrak{su}(2)$ for effective Hamiltonians of topological fermions under laser pumping, we provide a comprehensive understanding of chiral Floquet-engineered transient change of electronic structures in CoSi, which could be detected by the Mid-IR pumping and THz Kerr or Faraday probe spectroscopy measurements in an ultrafast time scale[57].
@@ 0.1 $\mathbf{k}\cdot\mathbf{p}$ model analysis
Fig. 2 $\bm{|}$: Chiral light induced momentum shifts for topological fermions in the LH CoSi. a,b The Floquet band structures around $\Gamma$ and R points for the LH CoSi under the radiation of the LCPL without SOC. Blue and red dot lines in (a) and (b) are obtained from the Floquet effective $\mathbf{k}\cdot\mathbf{p}$ model and the black dashed lines are obtained from the diagonalization of the Floquet tight-binding Hamiltonian. The cut-off of the Floquet index is set as $n=\{-1,0,1\}$, and the Fermi levels are set as zero. In these calculations, the photon energy is 100 meV and the electric field intensity is $4.4\times 10^{7}$ V/m. c,d The projection of the bulk states around $\Gamma$ and R points onto the (001) side surface respectively, where $\bar{\Gamma}$ and $\rm{\bar{M}}$ points are their projected points. The energy levels in (c) and (d) are set at crossing points of the spin-1 excitation and double Weyl fermion, respectively.
@@
Fig. 3 $\bm{|}$: Surface states in the LH and RH CoSi under chiral Floquet engineering. a The surface Floquet electronic structures along high symmetry lines, which are projected to the (001) surface for the LH (left panel) and RH (right panel) CoSi without SOC under the LCPL radiation. b The evolution of the constant-energy contours for Floquet surface states in the LH (left panel) and RH (right panel) CoSi without SOC. Three energy cuts are marked in (a) and the mirror plane is a gray parallelogram.
@@ 0.2 Tight-binding Hamiltonian analysis
Fig. 4 $\bm{|}$: The setup schematic of the pump-probe Kerr and Faraday experiments on the CoSi thin film sample. The CPL (pink) are illuminated onto the CoSi thin film. The polarization directions of the reflected and transmitted light are denoted as purple and blue arrows, and the Kerr ($\theta_{K}$) and Faraday ($\theta_{F}$) angles of the linearly polarized probe pulse (orange) are measured as a function of pump-probe delay time.
@@ Discussion
In this work, we demonstrate a chiral Floquet engineering of topological fermions in the CoSi family. However, the experimental realization of Floquet engineering is still challenging and only limited to a few material candidates[34, 59, 35, 60, 61, 62, 63, 64, 65, 66]. One significant obstacle in the formation of Floquet states lies in the competition between the laser pumping and the decoherent scattering for the excited electrons[67]. To avoid the Floquet state’s collapse, the pumping laser period ($T=2\pi/\Omega$) should be shorter than the experimental scattering time. For a metallic system, such as the surface state of $\text{Bi}_{2}\text{Te}_{3}$, a previous study has shown that the emergency of Floquet-Bloch sidebands can be built up within a few ultrashort optical cycles[35]. Based on the transport measurement[68], the scattering time in the CoSi single crystal is estimated as around 131 fs (see the Supplementary Note 6), which provides a lower limit to the pumping laser frequency. Consequently, considering these constraints, we expect a Mid-IR pumping laser with photon energy around 100 meV and an electric field intensity as large as $4.4\times 10^{7}$ V/m can be used to excite the Floquet states in CoSi compounds. Additionally, time- and angle-resolved photoemission spectroscopy (TrARPES) is a powerful and direct experimental technique[69] to observe Floquet band structures, such as the detection of Dirac surface[34, 35] and bulk[65, 66] states under the Mid-IR laser pumping (see the Supplementary Note 7 for details). Therefore, we propose that TrARPES can be used to probe the light-dressed electronic structures of both bulk and surface states.

### v2

@@ Introduction
In this Letter, we report the chiral Floquet engineering on topological fermions in the CoSi family using the Floquet effective $\mathbf{k\cdot p}$ model from the perturbation theory, complemented by the Floquet tight-binding Hamiltonian based on the ab initio calculations as a benchmark. We show that under CPL pumping, gapless topological fermions with different topological charges undergo momentum shifts along opposite directions (see Fig. 1b) while preserving their topological properties. This phenomenon is distinct from the momentum shifts of Weyl points driven by the light-induced shear phonon mode in WTe2[55] and shifts of topological fermions under the linearly polarized light (LPL)[56, 57]. The sign and magnitude for light-induced momentum shifts at $\Gamma$ and R points are determined by the Floquet chirality index $\Xi_{k=\Gamma,\rm{R}}$, which depends on the light-matter coupling strength and the interplay among the three distinct chiralities, the crystal handedness $\eta$, the chirality $\chi_{k=\Gamma,\rm{R}}(\eta)$ of topological fermions, and the chirality $\gamma$ of CPL. Via analyzing the Lie algebra representations of $\mathfrak{su}(2)$ for effective Hamiltonians of topological fermions under laser pumping, we provide a comprehensive understanding of chiral Floquet-engineered transient change of electronic structures in CoSi, which could be detected by the Mid-IR pumping and THz Kerr or Faraday probe spectroscopy measurements in an ultrafast time scale[58].
@@ Discussion
In this work, we demonstrate a chiral Floquet engineering of topological fermions in the CoSi family. However, the experimental realization of Floquet engineering is still challenging and only limited to a few material candidates[34, 60, 35, 61, 62, 63, 64, 65, 66, 67]. One significant obstacle in the formation of Floquet states lies in the competition between the laser pumping and the decoherent scattering for the excited electrons[68]. To avoid the Floquet state’s collapse, the pumping laser period ($T=2\pi/\Omega$) should be shorter than the experimental scattering time. For a metallic system, such as the surface state of $\text{Bi}_{2}\text{Te}_{3}$, a previous study has shown that the emergency of Floquet-Bloch sidebands can be built up within a few ultrashort optical cycles[35]. Based on the transport measurement[69], the scattering time in the CoSi single crystal is estimated as around 131 fs (see the Supplementary Note 6), which provides a lower limit to the pumping laser frequency. Consequently, considering these constraints, we expect a Mid-IR pumping laser with photon energy around 100 meV and an electric field intensity as large as $4.4\times 10^{7}$ V/m can be used to excite the Floquet states in CoSi compounds. Additionally, time- and angle-resolved photoemission spectroscopy (TrARPES) is a powerful and direct experimental technique[70] to observe Floquet band structures, such as the detection of Dirac surface[34, 35] and bulk[66, 67] states under the Mid-IR laser pumping (see the Supplementary Note 7 for details). It should be noted that a tunable probe photon energy is essential for the detection process[16, 14], particularly to track the evolution of topological fermions along the $k_{z}$ direction when subjected to the Mid-IR laser pumping that propagates along the $z$ direction. Therefore, we propose that TrARPES can be used to probe the light-dressed electronic structures of both bulk and surface states.
@@ 0.4 Floquet Theory
The code used for this study is not publicly available but may be made available to qualified researchers on reasonable request from the corresponding author.
P.T. was supported by the National Natural Science Foundation of China (Grants No. 12234011 and No. 12374053). B.F. and W.D. acknowledge the support of the Basic Science Center Project of NSFC (Grant No. 52388201), Innovation Program for Quantum Science and Technology (Grant No. 2023ZD0300500), and the Beijing Advanced Innovation Center for Future Chip (ICFC). A.R. acknowledges support from the Cluster of Excellence ’CUI: Advanced Imaging of Matter’- EXC 2056 - project ID 390715994, and Grupos Consolidados (IT1453-22). A.R. also acknowledges support from the Max Planck-New York City Center for Non-Equilibrium Quantum Phenomena. The Flatiron Institute is a division of the Simons Foundation.
P.T., D.W., and A.R. conceived and supervised the research project. B.F. developed the theoretical description and performed all numerical calculations. All authors contributed to the discussions and commented on the manuscript.
The authors declare no competing interests.
0.5 References